# NCL++: Nested Collaborative Learning for Long-Tailed Visual Recognition

### v1

@@ 4.4 Component Analysis
Influence of data augmentations. Data augmentation is a common tool to improve performance. For example, previous works zhang2021bag ; cai2021ace ; zhong2021improving use Mixup zhang2018mixup and RandAugment cubuk2020randaugment to obtain richer feature representations. Our method follows PaCo cui2021parametric to employ RandAugment cubuk2020randaugment for experiments. As shown in Table 4, the performance is improved by about 3% to 5% when employing RandAugment for training. However, our high performance depends not entirely on RandAugment. When using Mixup for augmentation, our method achieves the performance of 51.37% and 53.38% for single expert and ensemble of them respectively. Even with the simple data augmentation (random crop, random horizontal flip), our method still performs very well, the ensemble model reaches an amazing performance of 51.57%. The results show the superiority of our method over state-of-the-art methods for both simple and complex data augmentation.
Table 4: Comparisons of training the network with simple augmentation (SimAug), RandAugment (RandAug) and Mixup. Experiments are conducted on CIFAR100-LT dataset with an IF of 100. ${}^{\dagger}$ indicates the ensemble performance is reported.
@@
Current SOTA ${}^{\dagger}$ wang2020long 52.4 hou2022batchformer ${}^{\dagger}$ cai2021ace
@@
${}^{\dagger}$ 51.57 56.33 53.38
Analysis on computation cost. We compare the computation cost between the proposed method and other state-of-the-art (SOTA) methods as shown in Table 5. For a fair comparison, we set two experts equally for both NCL and NCL++. Due to using multiple networks for collaborative learning, NCL and NCL++ have more computation costs in the training stage. In the test stage, the NCL and its improved version, NCL++, only uses a single network for evaluation and it contains less computation cost but with a higher accuracy compared with other SOTA methods, like NCM kang2019decoupling and TCL li2021trustworthy . Moreover, compared with NCL++, NCL contains more computations (with higher GFLOPs) while a lower performance accuracy (57.4% vs. 58.0%). As we know, the only difference between NCL and NCL++ is that the self-supervision module is replaced with the newly proposed IntraCL module. Obviously, by replacing the self-supervision part with IntraCL, the computation cost is largely reduced while the performance is improved.
@@
NCM kang2019decoupling 4.11 4.11 44.3
TLC li2021trustworthy (3 experts) 6.55 6.55 55.1
NCL li2022nested (3 experts) 17.03 $*{}^{*}$ 57.4
NCL++ (2 experts) 8.24 $*{}^{*}$ 58.0
@@
    $*{}^{*}$ evaluating the model with a single expert.
Fixed vs. learnable loss weights. In this paper, we fixedly set the loss weights $\lambda_{1}$ and $\lambda_{2}$ and select the appropriate values for them by a cross-validation (see Fig. 5). The loss weights also can be learn automatically as opposed to the fixed values. In this sub-section, we set loss weights $\lambda_{1}$ and $\lambda_{2}$ as learnable parameters and compare it with the fixed manner. Following the work zhang2020self , we normalize the two learnable values $\lambda_{1}$ and $\lambda_{2}$ by a softmax function before using them to aggregate the losses, which ensures that $\lambda_{1}$ and $\lambda_{2}$ are non-negative values and $\lambda_{1}+\lambda_{2}=1$. The experimental results are shown in Table 6. As we can see, compared to the fixed weights, using the learnable loss weights can obtain a higher accuracy on the train set but receive a lower accuracy on the test set. It shows that the learnable loss weights may lead to overfitting and setting the parameters as fixed values may achieve better generalizations.
@@
[…]
@@ 4.5 Discussion and Further Analysis
t-SNE visualizations. We use the t-SNE van2008visualizing to visualize the features of the baseline method BSCE and the proposed NCL++. Compared to the baseline BSCE, the two-dimensional t-SNE map of the proposed NCL++ seems to be better clustered as shown in Fig. 10, (e.g., the red color). It shows that the proposed NCL++ generate more compact feature representation than the baseline BSCE, where the features of the same category extracted by the network stay closer. The nicely clustered and compact features indicate the model are well-trained for long-tailed visual recognition, which can achieve better classification and recognition performance.
Experiments on balanced datasets. Although NCL++ is proposed based on the phenomenon of long-tailed datasets, except for the re-balanced part, the rest components of NCL++ can still be applied to classification tasks on balanced datasets, such as CIFAR krizhevsky2009learning . We conduct experiments on CIFAR100 as shown in Table 8. Cross Entropy (denoted as ’CE’) is utilized as the baseline method and the result of the single expert is reported. As we can see from the results, the proposed NCL++ achieves considerable improvement over the baseline. The performance is improved by 2.61% with ResNet-10 and 3.20% with ResNet-32. It is undeniable that even in balanced datasets there are still confusing categories and the uncertainty of model training will still exist. In these scenarios, NCL++ also shows its advantages.
@@
Hybrid-PSC wang2021contrastive 45.0
PaCo cui2021parametric 52.0
NCL li2022nested 53.3
@@
Comparisons to contrastive learning methods. There are some similarities between our method and contrastive learning methods. For contrastive learning, it aims to narrow the distance between positive pairs (augment copies from the same image) by contrastive learning losses. For our NCL++, it employs the KL loss to minimize the distance between positive sample pairs. Specifically, the definition of the positive sample pairs are two folds: one is the different augment copies from the same image in our IntraCL, and the other is the different predictions of the same image from different networks in InterCL. However, there are also some differences among them. For example, contrastive learning is an unsupervised method while our method is a supervised method. It is precisely because of supervised learning where each sample has a certain label, we further propose the nested feature learning to increase the network’s ability to distinguish the sample from the negative hard categories. We have compared the performance of our method and other contrastive learning based methods (including Hybrid-PSC wang2021contrastive and PaCo cui2021parametric ) as shown in Table 10. As we can see, our method outperforms them by a considerable margin. Regarding the aspect of hard sample mining, Hybrid-PSC enhances training on the tail by reweighting, while PaCo enhances training on the tail by resampling. These approaches mine hard samples based on categories. While in our method, we proposed HCM to focus on each individual sample, extracting hard categories specific to that sample. Therefore, our method is still different from the hard example mining, and has not been proposed before. We further take a simple experiment to compare the performance of our IntraCL and a classic contrastive learning MoCo. Actually, when replacing the IntraCL with MoCo’s contrastive learning module, our NCL++ degenerates to the previous version NCL li2022nested . As shown in Table 10, our IntraCL could achieve better performance, which shows the advantages of the proposed method.

### v3

@@ 4.4 Component Analysis
Influence of data augmentations. Data augmentation is a common tool to improve performance. For example, previous works [30, 22, 72] use Mixup [81] and RandAugment [61] to obtain richer feature representations. Our method follows PaCo [66] to employ RandAugment [61] for experiments. As shown in Table 4, the performance is improved by about 3% to 5% when employing RandAugment for training. However, our high performance depends not entirely on RandAugment. When using Mixup for augmentation, our method achieves the performance of 51.37% and 53.38% for single expert and ensemble of them respectively. Even with the simple data augmentation (random crop, random horizontal flip), our method still performs very well, the ensemble model reaches an amazing performance of 51.57%. The results show the superiority of our method over state-of-the-art methods for both simple and complex data augmentation.
Table 4: Comparisons of training the network with simple augmentation (SimAug), RandAugment (RandAug) and Mixup. Experiments are conducted on CIFAR100-LT dataset with an IF of 100. indicates the ensemble performance is reported.
@@
Current SOTA 49.1 [20] 52.4 [78] 49.4 [22]
@@
BSCE+NCL++ 51.57 56.33 53.38
Analysis on computation cost. We compare the computation cost between the proposed method and other state-of-the-art (SOTA) methods as shown in Table 5. For a fair comparison, we set two experts equally for both NCL and NCL++. Due to using multiple networks for collaborative learning, NCL and NCL++ have more computation costs in the training stage. In the test stage, the NCL and its improved version, NCL++, only uses a single network for evaluation and it contains less computation cost but with a higher accuracy compared with other SOTA methods, like NCM [17] and TCL [44]. Moreover, compared with NCL++, NCL contains more computations (with higher GFLOPs) while a lower performance accuracy (57.4% vs. 58.0%). As we know, the only difference between NCL and NCL++ is that the self-supervision module is replaced with the newly proposed IntraCL module. Obviously, by replacing the self-supervision part with IntraCL, the computation cost is largely reduced while the performance is improved.
@@
NCM [17] 4.11 4.11 44.3
TLC [44] (3 experts) 6.55 6.55 55.1
NCL [67] (3 experts) 17.03 4.11 57.4
NCL++ (2 experts) 8.24 4.11 58.0
@@
    evaluating the model with a single expert.
Fixed vs. learnable loss weights. In this paper, we fixedly set the loss weights $\lambda_{1}$ and $\lambda_{2}$ and select the appropriate values for them by a cross-validation (see Fig. 5). The loss weights also can be learn automatically as opposed to the fixed values. In this sub-section, we set loss weights $\lambda_{1}$ and $\lambda_{2}$ as learnable parameters and compare it with the fixed manner. Following the work [54], we normalize the two learnable values $\lambda_{1}$ and $\lambda_{2}$ by a softmax function before using them to aggregate the losses, which ensures that $\lambda_{1}$ and $\lambda_{2}$ are non-negative values and $\lambda_{1}+\lambda_{2}=1$. The experimental results are shown in Table 6. As we can see, compared to the fixed weights, using the learnable loss weights can obtain a higher accuracy on the train set but receive a lower accuracy on the test set. It shows that the learnable loss weights may lead to overfitting and setting the parameters as fixed values may achieve better generalizations.
@@
Ablations studies on all components. In this sub-section, we perform detailed ablation studies for our NCL++ on CIFAR100-LT dataset, which is shown in Table 7. To conduct a comprehensive analysis, we evaluate the proposed components including Balanced Individual Learning (BIL), IntraCL, InterCL and the Nested Feature Learning on hard categories (denoted as ’NFLhard’). The ablation studies start from a naive method, which does not take any one of above components. To be specific, the naive method is indeed a plain network with taking the Cross-Entropy (CE) loss as the loss function. Compared with the naive method, the BIL improve the performance from 44.79% to 50.60%, which is a considerable improvement. This is because the naive method does not consider anything about the imbalanced data distribution. For the BIL, it dramatically increases the contributions of tail classes while suppressing that of head classes, which could greatly improve the performance by picking low hanging fruit. For other components like IntraCL, InterCL and NFLhard, all of them can steadily improve the performance. For example, based on BIL, the proposed InterCL and IntraCL improve the performance by 2.53% and 3.34%, respectively. When both InterCL and IntraCL are employed, this improvement can be enlarged to 3.76%. Then, the NFLhard also gains the improvements from 54.36% to 54.82%. Finally, we take the model ensemble of all experts (two experts are employed) to produce the final predictions, and the performance is further improved to 56.33%. The steadily performance improvements clearly show the effectiveness of the proposed NCL++.
Table 7: Ablation studies on CIFAR100-LT dataset with an IF of 100. ’BIL’, ’InterCL’ and ’IntraCL’ represent the individual learning or collaborative learning of the global view on all categories. For ’NFLhard’, it means adding the nested feature learning of the partial view on hard categories. ’Ensemble’ means that the final predictions are obtained by averaging the results on multiple experts. In the proposed NCL++, two experts are employed.
BIL InterCL IntraCL NFLhard Ensemble Acc.
@@ 4.5 Discussion and Further Analysis
t-SNE visualizations. We use the t-SNE [82] to visualize the features of the baseline method BSCE and the proposed NCL++. Compared to the baseline BSCE, the two-dimensional t-SNE map of the proposed NCL++ seems to be better clustered as shown in Fig. 10, (e.g., the red color). It shows that the proposed NCL++ generate more compact feature representation than the baseline BSCE, where the features of the same category extracted by the network stay closer. The nicely clustered and compact features indicate the model are well-trained for long-tailed visual recognition, which can achieve better classification and recognition performance.
Experiments on balanced datasets. Although NCL++ is proposed based on the phenomenon of long-tailed datasets, except for the re-balanced part, the rest components of NCL++ can still be applied to classification tasks on balanced datasets, such as CIFAR [64]. We conduct experiments on CIFAR100 as shown in Table 8. Cross Entropy (denoted as ’CE’) is utilized as the baseline method and the result of the single expert is reported. As we can see from the results, the proposed NCL++ achieves considerable improvement over the baseline. The performance is improved by 2.61% with ResNet-10 and 3.20% with ResNet-32. It is undeniable that even in balanced datasets there are still confusing categories and the uncertainty of model training will still exist. In these scenarios, NCL++ also shows its advantages.
@@
Hybrid-PSC [73] 45.0
PaCo [66] 52.0
NCL [67] 53.3
@@
Comparisons to contrastive learning methods. There are some similarities between our method and contrastive learning methods. For contrastive learning, it aims to narrow the distance between positive pairs (augment copies from the same image) by contrastive learning losses. For our NCL++, it employs the KL loss to minimize the distance between positive sample pairs. Specifically, the definition of the positive sample pairs are two folds: one is the different augment copies from the same image in our IntraCL, and the other is the different predictions of the same image from different networks in InterCL. However, there are also some differences among them. For example, contrastive learning is an unsupervised method while our method is a supervised method. It is precisely because of supervised learning where each sample has a certain label, we further propose the nested feature learning to increase the network’s ability to distinguish the sample from the negative hard categories. We have compared the performance of our method and other contrastive learning based methods (including Hybrid-PSC [73] and PaCo [66]) as shown in Table 10. As we can see, our method outperforms them by a considerable margin. Regarding the aspect of hard sample mining, Hybrid-PSC enhances training on the tail by reweighting, while PaCo enhances training on the tail by resampling. These approaches mine hard samples based on categories. While in our method, we proposed HCM to focus on each individual sample, extracting hard categories specific to that sample. Therefore, our method is still different from the hard example mining, and has not been proposed before. We further take a simple experiment to compare the performance of our IntraCL and a classic contrastive learning MoCo. Actually, when replacing the IntraCL with MoCo’s contrastive learning module, our NCL++ degenerates to the previous version NCL [67]. As shown in Table 10, our IntraCL could achieve better performance, which shows the advantages of the proposed method.